40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15222     Entry time: Mon Feb 24 08:36:32 2020
Author: Chub 
Type: Update 
Category: General 
Subject: HVAC repair 

The HVAC people replaced a valve and repaired the pneumatic plumbing on the roof air handler.  Temperature has been stable during the day since Thursday.  If anyone is in the control room during the evening, please make a note of the temperature.

Chub

ELOG V3.1.3-